# Experimental validation of a linear momentum and bluff-body model for high-blockage cross-flow turbine arrays

### v1

@@ Abstract
The performance and near-wake characteristics of a turbine in a confined flow depend on the blockage ratio, defined as the ratio of the turbine projected area to the channel cross-sectional area. While blockage is understood to increase the power coefficient for turbine “fences” spanning a channel, most investigations at the upper range of practically-achievable blockage ratios have been theoretical or numerical in nature. Furthermore, while linear momentum actuator disk theory is frequently used to model turbines in confined flows, as confinement increases, the ability of this idealized model to describe performance and flow fields has not been established. In this work, the performance and near-wake flow field of a pair of cross-flow turbines are experimentally evaluated at blockage ratios from $30\%$ to $55\%$. The fluid velocity measured in the bypass region is found to be well-predicted by the open-channel linear momentum model developed by Houlsby et al. (2008), while the wake velocity is not. Additionally, self-similar power and thrust coefficients are identified across this range of blockage ratios when array performance is scaled by the modeled bypass velocity following Whelan et al.’s (2009) adaptation of the bluff-body theory of Maskell (1963). This result demonstrates that, despite multiple non-idealities, relatively simple models can quantitatively describe highly confined turbines. From this, an analytical method for predicting array performance as a function of blockage is presented. Overall, this work illustrates turbine performance at relatively high confinement and demonstrates the suitability of analytical models to predict and interpret their hydrodynamics.
keywords:
1 Introduction
@@
Flowing water in river and tidal channels is a promising source of renewable power. Individual turbines and arrays deployed in these channels can convert the energy in the fluid into mechanical power, and the efficiency of such turbines is influenced by how much of the channel cross-section they occupy. The characteristic scale of a row of turbines in a channel is represented by the blockage ratio, defined as the ratio between the turbine projected area and the channel cross-sectional area:
$\beta=\frac{\mathrm{A_{turbines}}}{A_{\mathrm{channel}}}.$ (1)
@@
As the blockage ratio increases, the turbines present greater resistance to the oncoming flow. For a constant volumetric flow rate, this resistance, combined with confinement from the channel boundaries, increases flow through the turbine rotors relative to unconfined conditions. As a consequence, as confinement increases, so does turbine efficiency (Garrett & Cummins, 2007; Whelan et al., 2009; Nishino & Willden, 2012b; Vogel et al., 2016). While blockage augments the performance of all turbine designs, cross-flow turbines—which have an axis of rotation perpendicular to the direction of inflow—are particularly well-suited to exploit these physics since their rectangular projected area allows them to more easily achieve high blockage than circular axial-flow turbines.
@@
Garrett & Cummins (2007) were the first to establish the theoretical basis for turbine performance in confined flows by applying one-dimensional linear momentum actuator disk theory (LMADT) to a turbine in a tidal channel. For relatively low Froude numbers, they showed that the maximum efficiency in confined flow was augmented by a factor of $(1-\beta)^{-2}$. Subsequent studies further generalized this analytical model. Houlsby et al. (2008) and Whelan et al. (2009) relaxed the assumption of low Froude number by allowing the free surface to deform in response to momentum extraction. Subsequently, Nishino & Willden (2012b, 2013) introduced a two-scale LMADT model for turbine arrays that considered confinement effects from both the channel boundaries (“global” blockage) and the presence of neighboring turbines within the array (“local” blockage). Vogel et al. (2016) later combined these approaches into a single model that incorporated both free surface deformation and inter-turbine proximity effects, and Dehtyriov et al. (2021) extended the two-scale model of Nishino & Willden (2012b) to an arbitrary number of scales (i.e., arrays of turbine sub-arrays) using a fractal-like approach.
@@
A summary of prior experimental and numerical work evaluating turbine performance at multiple blockage ratios is provided in table 1. Across these studies, performance results are generally consistent and aligned with theory: for all types of turbines, as blockage increases, so do the thrust loading on the turbine, maximum efficiency, and optimal operating speed (i.e., tip-speed ratio or reduced frequency). Studies that examine the flow field near confined turbines (Battisti et al., 2011; Nishino & Willden, 2012a; McTavish et al., 2014; Dossena et al., 2015; Schluntz & Willden, 2015; Sarlak et al., 2016; Ross & Polagye, 2020b) find that an increase in blockage accelerates the flow through and around the turbine as predicted by theory, as well as narrows the turbine wake. Gauthier et al. (2016) and Kinsey & Dumas (2017) examined how blockage effects change with confinement asymmetry (i.e., different channel widths and heights at constant $\beta$) for various turbine types, and found turbine performance to be invariant to confinement asymmetry for width-to-height and height-to-width ratios approximately less than $3$. The interplay between rotor geometry and blockage effects has also been explored for cross-flow turbines by Goude & Ågren (2014) and for axial-flow turbines by Schluntz & Willden (2015) and Abutunis & Menta (2022). These studies found that higher-solidity rotors benefit more from blockage – a consequence of how their higher thrust coefficients couple with channel-level energetics.
Several of the studies in table 1 also explore analytical blockage corrections—which adjust observed turbine performance in confined flow to predict the performance of the same turbine in unconfined flow—and investigate their efficacy experimentally (Whelan et al., 2009; Battisti et al., 2011; Chen & Liou, 2011; Ross & Altman, 2011; Dossena et al., 2015; Jeong et al., 2018; Ross & Polagye, 2020a) or numerically (Gauthier et al., 2016; Kinsey & Dumas, 2017; Zilic de Arcos et al., 2020; Zhang et al., 2023). LMADT models are often the physical basis for blockage corrections (e.g., Barnsley & Wellicome, 1990; Whelan et al., 2009; Ross & Polagye, 2020a; Steiros et al., 2022; Dehtyriov et al., 2023). Such methods generally proceed from Glauert (1935) (originally introduced as a blockage correction for airplane propellers) by assuming that the thrust on a turbine depends on the incident freestream velocity, with LMADT employed to relate the confined and unconfined freestream velocities. In contrast, Whelan et al. (2009) found that an LMADT-based blockage correction inspired by the bluff-body theory of Maskell (1963) could be more effective for predicting the unconfined thrust coefficients of highly-loaded turbines. Unlike Glauert’s theory, a Maskell-inspired correction assumes that the thrust on a turbine depends on the velocity of the accelerated flow that bypasses it. However, none of these studies have demonstrated that the underlying LMADT models are actually descriptive of the flow field around confined turbines.
@@
Author Turbine Type $\beta$ [%] Author Turbine Type $\beta$ [%]
Takamatsu et al. (1985) CFT 75, 95 Nishino & Willden (2012a) Actuator disk 3.1 - 50.3
Whelan et al. (2009) AFT 5, 64 Consul et al. (2013) CFT 12.5, 25, 50
McAdam et al. (2013a, b) CFT 47, 59 Goude & Ågren (2014) CFT 0, 12.5, 25, 50
Battisti et al. (2011) CFT 2.8, 10 Kolekar & Banerjee (2015) AFT 4.2, 8.2, 11, 17, 20, 33, 42
Chen & Liou (2011) AFT 10.2, 20.2, 28.3 Schluntz & Willden (2015) AFT 0, 7.9, 19.6, 31.4
Ross & Altman (2011) CFT (Savonius) 2, 3.5, 8 Gauthier et al. (2016) Oscillating foil 0.2 - 60.5
Birjandi et al. (2013) CFT 24.5 - 49.2 Sarlak et al. (2016) AFT 2, 5, 9, 20
McTavish et al. (2014) AFT 6.3, 9.9, 14.3, 19.4, 25.4 Kinsey & Dumas (2017) AFT 0, 10, 20, 50, 60
Gaurier et al. (2015) AFT 1.2, 3.3, 4.8 CFT 0, 13, 26, 51
Ryi et al. (2015) AFT 8.1, 18.0, 48.1 Badshah et al. (2019) AFT 2 - 19
Dossena et al. (2015) CFT 2.8, 10 Gauvin-Tremblay & Dumas (2020) CFT 5, 20, 40
Jeong et al. (2018) CFT 3.5, 13.4, 24.7 Zilic de Arcos et al. (2020) AFT 1, 5, 10, 20, 40
Ross & Polagye (2020a) AFT 2, 35 Abutunis & Menta (2022) AFT 4.2, 20.8, 41.6, 62.5
CFT 3, 36 Zhang et al. (2023) AFT 2, 2.8, 4.4, 7.9, 17.7
Ross & Polagye (2020b) CFT 14, 36 CFT (Savonius) 2.5, 3.6, 5.6, 10, 22.5
@@
Although the studies in table 1 may appear to provide a comprehensive understanding of how confinement affects turbine performance, the upper end of blockage ratios that are achievable in a realistic channel ($\beta=30\%-60\%$) have been primarily explored through numerical simulation (e.g., Consul et al., 2013; Kinsey & Dumas, 2017; Abutunis & Menta, 2022). While the general trends in turbine performance and wake characteristics observed in simulation at these higher blockages are in agreement with theory, these have not been validated, in part due to a lack of equivalent experimental data. Only a handful of the experimental studies in table 1 consider blockage ratios greater than $30\%$, and of this subset, only Ross & Polagye (2020b) characterize the turbine wake. Furthermore, experimental studies that do explore the upper-end of blockage ratios consider only a sparse set of blockages (likely due to the difficulty in achieving a wide range of blockages while maintaining dynamic similarity (Hunt & Polagye, 2023)) which limits the ability to identify how performance trends evolve as a function of blockage. In addition, meaningful cross-comparisons between experiments are complicated by differences in rotor geometry, which can strongly influence performance (Schluntz & Willden, 2015; Hunt et al., 2024), and several studies convolve changes in blockage with changes in the Reynolds and Froude numbers (e.g., McAdam et al., 2013a, b; Birjandi et al., 2013). Given the lack of high-resolution experimental data at higher blockages, the computational methods used in numerical studies are often validated only at lower blockage (Consul et al., 2013; Kolekar & Banerjee, 2015; Gauthier et al., 2016; Kinsey & Dumas, 2017)—or not at all—before being used to draw conclusions at higher blockage. Consequently, although the general effects of blockage on turbine performance are well-known, there is residual uncertainty as to how these effects evolve at higher confinement. For these reasons, a high-resolution experimental examination of turbine performance and the associated near-wake flow field across the upper end of practically-achievable blockage ratios is warranted. In doing so, the accuracy of LMADT models for describing the near-wake flow field may also be evaluated, providing insight into the mechanics of blockage corrections.
@@
Figure 1: Outline of the procedure in Sections 3 and 4 to interpret high blockage cross-flow turbine dynamics using experimental performance data, flow fields, and linear momentum theory. Key methodologies that are adapted from prior work are highlighted in gray boxes.
@@
The objective of this work is to develop a more comprehensive understanding of performance and flow fields around a cross-flow turbine array at higher blockages. Such information can be used to understand the relevance of theory to non-ideal arrays that arguably violate multiple assumptions that underpin LMADT (e.g., treating the turbines as porous, non-rotating disks). Similarly, these experimental results provide a high-resolution data set for future validation of simulations at high confinement. To this end, a symmetric two-turbine array is tested in a recirculating water channel at blockage ratios ranging from $30\%$ to $55\%$, and particle image velocimetry (PIV) is used to measure the flow field in the near wake. In these experiments, the Reynolds and Froude numbers are controlled to avoid convolving changes in blockage with other relevant non-dimensional parameters (Ross & Polagye, 2022). Section 2 describes the experimental methods used to characterize array performance and flow fields, with trends in performance, near-wake velocities, and free surface deformation discussed in Section 3. In Section 4, the open-channel LMADT model developed by Houlsby et al. (2008) is applied to the experimental performance data to model the wake velocity, bypass velocity, and free surface deformation in the vicinity of the array. The analytically-modeled velocities are compared to measurements and are used to interpret high-blockage turbine hydrodynamics following Whelan et al.’s (2009) adaptation of the bluff-body theory of Maskell (1963). The procedure followed in Sections 3 and 4 is graphically outlined in figure 1 and methodologies that are adapted from prior work are highlighted. Based on the characteristic dynamics observed, a method for forecasting confined turbine performance based on linear momentum theory and Maskell’s (1963) bluff body theory (i.e., the reverse of the procedure in figure 1) is presented in Section 5.
2 Experimental Methods
2.1 Cross-Flow Turbine Test Setup
@@
Experiments are conducted in the Alice C. Tyler recirculating water flume at the University of Washington. The flume has a test section that is 0.76 m wide and 4.88 m long, and can accommodate water depths up to 0.60 m and flow speeds up to ${\sim}\!1.1$ m/s. The flume is equipped with both a heater and chiller for temperature control, and can maintain water temperatures between $10$ ${}^{\circ}\mathrm{C}$ and $40$ ${}^{\circ}\mathrm{C}$.
@@
The laboratory-scale array consists of two identical straight-bladed cross-flow turbines. The use of a pair of turbines, rather than an individual turbine, is motivated by the application of the research to understand the potential for confinement to reduce energy costs for turbine arrays. The rotors are two-bladed, consisting of a NACA 0018 profile with a 0.0742 m chord length ($c$) and mounted at a $-6^{\circ}$ (i.e., toe-out) preset pitch angle as referenced from the quarter chord. The blade span ($H$) is 0.215 m and the turbine diameter ($D$) is 0.315 m, measured at the outermost point swept by the blades. The blades are attached to the central driveshaft of each rotor using thin, hydrodynamic blade-end struts (NACA 0008 profile, 0.0742 m chord length). This arrangement results in an aspect ratio ($H/D$) of 0.68, chord-to-radius ratio ($c/R$) of 0.47, and solidity (product of the blade count and chord length relative to rotor circumference: $Nc/\pi D$, where $N$ is the blade count) of 0.15.
@@
The two rotors, designated as “Turbine A” and “Turbine B”, are integrated into the experimental setup shown in figure 3, which consists of two identical test-rigs. The top of each turbine’s central shaft is connected by a flexible shaft coupling (Zero-Max SC040R) to a servomotor (Yaskawa SGMCS-05BC341) which regulates the rotation rate of the turbine. This control strategy yields similar time-average performance (Polagye et al., 2019) and wake characteristics (Araya & Dabiri, 2015) to applying a constant oppositional torque and reduces cycle-to-cycle variability . The angular position of each turbine is measured via the servomotor encoder ($2^{16}$ counts per revolution), from which the angular velocity is estimated. The bottom of each turbine’s central shaft sits in a bearing. The net forces and torques on each turbine are measured by a pair of 6-axis load cells: an upper load cell (ATI Mini45-IP65) connected to the servomotor through a thermal isolation assembly and fixed to a rigid crossbeam, and a lower load cell (ATI Mini45-IP68) mounted to the bottom bearing and fixed to the bottom of the flume via a suction plate. Measurements from the load cells and servomotor encoders for both turbines are acquired synchronously at 1000 Hz in MATLAB using two National Instruments PCIe-6353 DAQs.
The freestream velocity, $U_{\infty}$, is measured using an acoustic Doppler velocimeter (Nortek Vectrino Profiler) sampling at 16 Hz. The velocimeter sampled a single cell positioned laterally in the center of the flume, vertically at the array midplane, and $5D$ upstream of the turbines’ axes of rotation. Velocity measurements are despiked using the method of Goring & Nikora (2002). The water depth far upstream of the array ($5.8D$ upstream of the array centerline, centered laterally in the channel) as well as in the vicinity of the array ($0.5D$, $1.0D$, and $2.0D$ directly upstream and downstream of Turbine B) is measured using seven synchronized ultrasonic free surface transducers sampling at 0.5 Hz. As the array is symmetric about the channel centerline (figure 3) the streamwise free surface profile in the vicinity of Turbine B is assumed to be representative of that in the vicinity of Turbine A. The water temperature is measured using a temperature probe (Omega Ultra-Precise RTD) and maintained within ±$\pm 0.1^{\circ}\mathrm{C}$ of the target value during each experiment.
2.2 Non-Dimensional Parameters

### v2

@@ Abstract
In a confined flow, the performance of a turbine and its near-wake fluid dynamics depend on the blockage ratio, defined as the ratio of the turbine projected area to the channel cross-sectional area. While blockage is understood to increase the power coefficient for turbine “fences” spanning a channel, most investigations at the upper range of practically-achievable blockage ratios have been theoretical or numerical in nature. Furthermore, while linear momentum actuator disk theory is frequently used to model turbines in confined flows, as confinement increases, the ability of this idealized model to describe performance and flow fields has not been established. In this work, the performance and near-wake flow field of a pair of cross-flow turbines are experimentally evaluated at blockage ratios from $30\%$ to $55\%$. The fluid velocity measured in the bypass region is found to be well-predicted by the open-channel linear momentum model developed by Houlsby et al. [1], while the wake velocity is not. Additionally, self-similar power and thrust coefficients are identified across this range of blockage ratios when array performance is scaled by the modeled bypass velocity following Whelan et al.’s [2] adaptation of the bluff-body theory of Maskell [3]. This result demonstrates that, despite multiple non-idealities, relatively simple models can quantitatively describe highly confined turbines. From this, an analytical method for predicting array performance as a function of blockage is presented. Overall, this work illustrates turbine performance at relatively high confinement and demonstrates the suitability of analytical models for predicting and interpreting their hydrodynamics.
Suggested keywords
preprint: es2025mar20_790
I Introduction
@@
Flowing water in river and tidal channels is a promising source of renewable power. The efficiency of individual turbines or arrays deployed in these channels is influenced by how much of the channel cross-section they occupy. For a row of turbines, this is represented by the blockage ratio, defined as the ratio between the turbine projected area and the channel cross-sectional area:
$\beta=\frac{A_{\mathrm{turbines}}}{A_{\mathrm{channel}}}.$ (1)
@@
As the blockage ratio increases, the turbines present greater flow resistance, but confinement from the channel boundaries leads to increased flow through the turbines relative to unconfined conditions. Consequently, as confinement increases, so does turbine efficiency [4, 2, 5, 6]. While blockage augments the performance of all turbine designs, cross-flow turbines—which have an axis of rotation perpendicular to the direction of inflow—are particularly well-suited to exploit these physics since their rectangular projected area matches the profile of natural and constructed channels.
@@
Garrett and Cummins [4] were the first to establish the theoretical basis for turbine performance in confined flows by applying one-dimensional linear momentum actuator disk theory (LMADT). For relatively low Froude numbers, they showed that the maximum efficiency in confined flow was augmented by a factor of $(1-\beta)^{-2}$. Subsequent studies by Houlsby et al. [1] and Whelan et al. [2] relaxed the assumption of low Froude number by allowing the free surface to deform in response to momentum extraction. In parallel, Nishino and Willden [5, 7] expanded the conceptual framework with a two-scale LMADT model for turbine arrays that considered confinement effects from both the channel boundaries (“global” blockage) and the presence of neighboring turbines within the array (“local” blockage). Vogel et al. [6] later combined these approaches into a single model that incorporated both free surface deformation and inter-turbine proximity effects, and Dehtyriov et al. [8] extended the two-scale model of Nishino and Willden [5] to an arbitrary number of scales (i.e., arrays of turbine sub-arrays) using a fractal-like approach.
@@
Blockage effects have been the focus of multiple experimental and numerical studies (Table 1). Across these, turbine dynamics are generally consistent and aligned with theory: for all types of turbines, as blockage increases, so do the rotor thrust loading, maximum efficiency, and optimal operating speed (i.e., tip-speed ratio or reduced frequency). Studies that examine the flow field near confined turbines [9, 10, 11, 12, 13, 14, 15] find that an increase in blockage accelerates the flow through and around the turbine as predicted by theory, as well as narrows the turbine wake. Gauthier et al. [16] and Kinsey and Dumas [17] examined how blockage effects change for various turbine types in asymmetric confinement (i.e., different combinations of lateral blockage and vertical blockage for a given $\beta$). The authors parameterized confinement asymmetry as the ratio between lateral blockage and vertical blockage (or its reciprocal, if a larger value), and found turbine performance to be invariant for values of confinement asymmetry less than three. The interplay between rotor geometry and blockage effects has also been explored for cross-flow turbines by Goude and Ågren [18] and for axial-flow turbines by Schluntz and Willden [13] and Abutunis and Menta [19]. These studies found that higher-solidity rotors benefit more from blockage – a consequence of how their higher thrust coefficients couple with channel-level energetics.
Several groups have investigated analytical blockage corrections—which adjust observed turbine performance in confined flow to predict the performance of the same turbine in unconfined flow—and evaluated those corrections through experiments [2, 9, 20, 21, 12, 22, 23, 24] or simulations [16, 17, 25, 24, 26]. Such methods generally proceed from Glauert [27] (who originally proposed a blockage correction for airplane propellers) by assuming that there exists some freestream velocity in unconfined flow that yields the same thrust and velocity through the rotor as in the confined condition. This unconfined freestream velocity is found through an analytical model, often LMADT [e.g., 28, 2, 23, 29]. However, blockage corrections based on LMDAT can yield non-physical results for “highly-loaded” turbines since sufficiently high thrust coefficients necessitate reversed flow in the turbine wake [30, 31]. This limitation has prompted investigation into blockage corrections that are applicable to a broader range of turbine solidities and operating conditions. For example, Whelan et al. [2] proposed a LMADT-based correction inspired by the bluff-body theory of Maskell [3], which, unlike Glauert’s theory, assumes that the thrust on a turbine depends on the velocity of the accelerated flow that bypasses it (i.e., the rotor thrust and bypass velocity are unchanged between corresponding confined and unconfined conditions). Whelan et al. found that such a correction more accurately predicted unconfined thrust coefficients for highly-loaded turbines. Recently, Steiros et al. [24] developed an alternative blockage correction for high-solidity turbines that proceeds from Glauert’s theory, but replaces LMADT with a two-dimensional potential flow model following Taylor’s [32] theory for drag on a porous plate, with adjustments for mass conservation [33] and wake pressure [30]. Steiros et al. [24] demonstrated that this approach overcame the limitations of LMADT for higher thrust coefficients, while maintaining good agreement with the more conventional method at lower confinements and solidities. However, when validating their model with drag measurements on porous plates, Steiros et al. [24] observed increasing disagreement between their theory and experiments with increasing confinement and solidity, which leaves open the question of whether a bluff-body correction could be more accurate in this regime.
@@
Author Turbine Type $\beta$ tested [%] Author Turbine Type $\beta$ tested [%]
Takamatsu et al. [34] CFT 75, 95 Nishino and Willden [10] Actuator disk 3.1 - 50.3
Whelan et al. [2] AFT 5, 64 Consul et al. [35] CFT 12.5, 25, 50
McAdam et al. [36, 37] CFT 47, 59 Goude and Ågren [18] CFT 0, 12.5, 25, 50
Battisti et al. [9] CFT 2.8, 10 Kolekar and Banerjee [38] AFT 4.2, 8.2, 11, 17, 20, 33, 42
Chen and Liou [20] AFT 10.2, 20.2, 28.3 Schluntz and Willden [13] AFT 0, 7.9, 19.6, 31.4
Ross and Altman [21] CFT (Savonius) 2, 3.5, 8 Gauthier et al. [16] Oscillating foil 0.2 - 60.5
Birjandi et al. [39] CFT 24.5 - 49.2 Sarlak et al. [14] AFT 2, 5, 9, 20
McTavish et al. [11] AFT 6.3, 9.9, 14.3, 19.4, 25.4 Kinsey and Dumas [17] AFT 0, 10, 20, 50, 60
Gaurier et al. [40] AFT 1.2, 3.3, 4.8 CFT 0, 13, 26, 51
Ryi et al. [41] AFT 8.1, 18.0, 48.1 Badshah et al. [42] AFT 2 - 19
Dossena et al. [12] CFT 2.8, 10 Gauvin-Tremblay and Dumas [43] CFT 5, 20, 40
Jeong et al. [22] CFT 3.5, 13.4, 24.7 Zilic de Arcos et al. [25] AFT 1, 5, 10, 20, 40
Ross and Polagye [23] AFT 2, 35 Abutunis and Menta [19] AFT 4.2, 20.8, 41.6, 62.5
CFT 3, 36 Steiros et al. [24] AFT 3.1, 19.6
Ross and Polagye [15] CFT 14, 36 Zhang et al. [26] AFT 2, 2.8, 4.4, 7.9, 17.7
CFT (Savonius) 2.5, 3.6, 5.6, 10, 22.5
@@
Although the studies in Table 1 may appear to comprehensively explore the effects of confinement on turbine performance, the upper end of blockage ratios that are achievable in a realistic river or tidal channel ($\beta=30\%-60\%$) have been primarily explored through numerical simulation. While the general trends in turbine performance and wake characteristics observed in these simulations are in agreement with theory, the computational methods employed are often validated only at lower blockage [35, 38, 16, 17]—or not at all—before investigating higher blockages. This is likely due to a lack of equivalent experimental data at high blockage ratios; only a handful of the experimental studies have considered blockage ratios greater than $30\%$. Experimental studies that do explore the upper-end of blockage ratios consider only a sparse set of $\beta$, likely due to difficulty of maintaining dynamic similarity across a wider range of blockage ratios [44]. Furthermore, meaningful cross-comparisons between experimental studies are complicated by differences in rotor geometry, and several studies simultaneously vary blockage along with Reynolds and Froude numbers [e.g., 36, 37, 39]. Consequently, although the general effects of blockage on turbine performance are well-known, there remains uncertainty as to how these effects evolve at higher confinement. Additionally, despite the widespread application of analytical blockage corrections in prior work, their suitability for higher confinements has not been fully established, nor has their ability to describe the flow field around confined turbines been thorough explored.
[…]
@@
Figure 1: Outline of the procedure in Sections III and IV to interpret high blockage cross-flow turbine dynamics using 1 experimental performance data, 1 flow fields, and 1 linear momentum theory, resulting in 1 self-similar array performance across blockage ratios. Key methodologies that are adapted from prior work are highlighted in gray boxes.
@@
The objective of this work is to develop a more comprehensive understanding of performance and flow fields around a cross-flow turbine array at higher blockages than previously considered in most experiments. Such information can be used to understand the relevance of theory to non-ideal arrays that arguably violate multiple assumptions underpinning analytical models (e.g., treating the turbines as porous, non-rotating disks in LMADT). To this end, a symmetric two-turbine array is tested in a recirculating water channel at blockage ratios ranging from $30\%$ to $55\%$, and flow fields in the near wake are characterized using particle image velocimetry (PIV). The Reynolds and Froude numbers are controlled to avoid convolving changes in blockage with other relevant non-dimensional parameters [45]. Section II describes the experimental methods used to characterize array performance and flow fields, with resulting performance, near-wake velocities, and free surface deformation discussed in Section III. In Section IV, the open-channel LMADT model developed by Houlsby et al. [1] is applied to the experimental performance data to model the wake velocity, bypass velocity, and free surface deformation in the vicinity of the array. The analytically-modeled velocities are compared to measurements and are used to interpret high-blockage turbine hydrodynamics following Whelan et al.’s [2] adaptation of the bluff-body theory of Maskell [3]. The procedure followed in Sections III and IV is graphically outlined in Figure 1 and methodologies that are adapted from prior work are highlighted. Based on the characteristic dynamics observed, a method for forecasting confined turbine performance based on LMADT and Maskell’s [3] bluff body theory (i.e., the reverse of the procedure in Figure 1) is presented in Section V.
II Experimental Methods
II.1 Cross-Flow Turbine Test Setup
@@
Experiments are conducted in the Alice C. Tyler recirculating water flume at the University of Washington. The flume has a test section that is 4.88 m long with a width ($w$) of 0.76 m, and can accommodate water depths up to 0.60 m and flow speeds up to ${\sim}\!1.1$ m/s. The flume is equipped with both a heater and chiller for temperature control, and can maintain water temperatures between $10$ ${}^{\circ}\mathrm{C}$ and $40$ ${}^{\circ}\mathrm{C}$.
@@
The laboratory-scale array consists of two identical straight-bladed cross-flow turbines. Experiments with a pair of turbines, rather than a single turbine, are motivated by the research application to confinement benefits for turbine arrays. The rotors are two-bladed, with each blade consisting of a NACA 0018 profile with a 0.0742 m chord length ($c$) mounted at a $-6^{\circ}$ (i.e., toe-out) preset pitch angle as referenced from the quarter chord. The blade span ($H$) is 0.215 m and the turbine diameter ($D$) is 0.315 m, measured at the outermost point swept by the blades. The blades are attached to the central driveshaft of each rotor using thin, hydrodynamic blade-end struts (NACA 0008 profile, 0.0742 m chord length). This arrangement results in an aspect ratio ($H/D$) of 0.68, chord-to-radius ratio ($c/R$) of 0.47, and solidity ($Nc/\pi D$, where $N$ is the blade count) of 0.15.
@@
The two rotors, designated as “Turbine A” and “Turbine B”, are integrated into the experimental setup shown in Figure 3, which consists of two identical test-rigs. The top of each turbine’s central shaft is connected by a flexible shaft coupling (Zero-Max SC040R) to a servomotor (Yaskawa SGMCS-05BC341) which regulates the rotation rate of the turbine. This control strategy yields similar time-average performance [46] and wake characteristics [47] to applying a constant oppositional torque and reduces cycle-to-cycle variability . The angular position of each turbine is measured via the servomotor encoder ($2^{16}$ counts per revolution), from which the angular velocity is estimated. The bottom of each turbine’s central shaft sits in a bearing. The net forces and torques on each turbine are measured by a pair of 6-axis load cells: an upper load cell (ATI Mini45-IP65) connected to the servomotor through a thermal isolation assembly and fixed to a rigid crossbeam, and a lower load cell (ATI Mini45-IP68) mounted to the bottom bearing and fixed to the bottom of the flume via a suction plate. Measurements from the load cells and servomotor encoders for both turbines are acquired synchronously at 1000 Hz in MATLAB using two National Instruments PCIe-6353 DAQs.
The freestream velocity, $U_{\infty}$, is measured using an acoustic Doppler velocimeter (Nortek Vectrino Profiler) sampling at 16 Hz. The velocimeter samples a single cell positioned laterally in the center of the flume, vertically at the turbine midspan, and $5D$ upstream of the turbines’ axes of rotation. Velocity measurements are despiked using the method of Goring and Nikora [48]. The turbulence intensity of the freestream flow measured by the ADV is ${\sim}2\%$ for all experiments. The water depth far upstream of the array ($5.8D$ upstream of the turbines’ axes of rotation, centered laterally in the channel) and depths in vicinity of the array ($0.5D$, $1.0D$, and $2.0D$ directly upstream and downstream of Turbine B) are measured using seven synchronized ultrasonic free surface transducers (Omega LVU32) sampling at 0.5 Hz. As the array is symmetric about the channel centerline (Figure 3) the streamwise free surface profile in the vicinity of Turbine B is assumed to be representative of that in the vicinity of Turbine A. The water temperature is measured using a temperature probe (Omega Ultra-Precise RTD) and maintained within ±$\pm 0.1^{\circ}\mathrm{C}$ of the target value during each experiment.
II.2 Non-Dimensional Parameters